OpenStudy (r_josh10):

A puppy weighed 27 lb. His weight increased by 233% by the time he was 1 year old. How much did the puppy weigh when he was 1 year old?

OpenStudy (r_josh10):

I need help with this question. Will medal and fan. Thank You.

563blackghost (563blackghost):

27 x .233 <--- %233 changed to a decimal = 6.291 + 27 = (33.291 lb)

OpenStudy (r_josh10):

ok thank You @563blackghost .

563blackghost (563blackghost):

hold on

563blackghost (563blackghost):

sorry i did the decimal wrong 27 x 2.33 = 62.91 +27 = 89.91